# Effects of Algorithmic Visibility on Conspiracy Communities: Reddit after Epstein’s ‘Suicide’

### v1

@@ Abstract.
This paper examines how algorithmic visibility shapes a large conspiracy community on Reddit after Jeffrey Epstein’s death. We ask whether homepage exposure changes who join r/conspiracy, how long they stay, and how they adapt linguistically, compared with users who arrive through organic discovery. Using a computational framework that combines toxicity scores, survival analysis, and lexical and semantic measures, the study shows that homepage visibility acts as a selection mechanism rather than a simple amplifier. Users who discover the community organically integrate more quickly into its linguistic and thematic norms and show more stable engagement over time. By contrast, users who arrive through visibility on the homepage remain semantically distant from core discourse and participate more briefly. Overall, algorithmic visibility reshapes audience size, community composition, and linguistic cohesion: newcomers who do not join organically have different incentives, integrate weakly, and leave quickly, which limits organic growth. In this high-risk setting, the observed behavioral and linguistic trajectories over five months do not match standard narratives in which incidental exposure to conspiracy content produces durable radicalization. These findings can inform the design of web platforms and recommendation systems that seek to curb harmful conspiracy exposure while supporting more responsible, transparent, and socially beneficial uses of algorithmic recommendations.
algorithmic visibility, online communities, conspiracy theories, Reddit, toxicity, user retention, linguistic integration, radicalization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Social networksccs: Human-centered computing Empirical studies in collaborative and social computing
1. Introduction
@@
Conspiracy theories are false or unverified narratives that allege secret, often sinister plots or events involving groups, organizations, or governments (Douglas and Sutton, 2023). They pose significant risks since they can change individuals’ perception of real events (Uscinski et al., 2017). Social media platforms provide especially fertile ground for their circulation, in part due to the emergence of fringe online communities (Rollo et al., 2022; Spohr, 2017). Such theories range from relatively benign conjectures to more harmful beliefs that amplify misinformation, heighten polarization, and potentially lead to serious real-world consequences.
@@
Online conspiracy communities are critical environments for understanding how digital platforms shape ideological discourse. Yet a crucial gap remains: the role of algorithmic visibility in determining who joins these communities, how long they stay, and whether they culturally integrate (Klein et al., 2019; Phadke et al., 2022; Venturini, 2022; Zeng et al., 2022; Corso et al., 2025b). Existing work on conspiracy engagement shows that users follow distinct pathways into and through conspiracy spaces and that linguistic adaptation and community ties play a central role in deeper involvement (Klein et al., 2019; Phadke et al., 2022). At the same time, research on platform interventions and deplatforming shows that conspiracy communities can be unusually resilient and mobile, rebuilding networks elsewhere even after bans (Monti et al., 2023). What remains unclear is how platform-level visibility itself, before any formal intervention, shapes who enters these communities and what kinds of membership emerge.
@@
Reddit offers a useful setting for such a research question (Proferes et al., 2021). It organizes content into topical subreddits and uses ranking and recommendation systems to decide which posts appear on the site’s homepage. When a subreddit reaches the homepage, it exposes its community to millions of users who are not subscribed to it and might not otherwise encounter its content. The subreddit r/conspiracy is Reddit’s largest conspiracy community, with more than two million members and a long record of discussions that range from historical events to contemporary politics. This combination of scale, public visibility, and stable infrastructure makes r/conspiracy an ideal case for studying how visibility events affect community composition, participation, and language (de Wildt and Aupers, 2024).
@@
The death of financier Jeffrey Epstein on August 10, 2019, created a sharp visibility shock (Schatto-Eckrodt et al., 2024). The peculiar circumstances of his death aligned fringe conspiracy discourse with mainstream skepticism and brought r/conspiracy and its 2 million members to the fore, reaching the platform’s homepage (noa, 2019; Ellis, [n. d.]). This visibility event changed the community’s reach without changing its formal rules or moderation structures, and it does so over a short, well-defined time window. It thus represents a perfect natural experiment.
@@
This paper examines how Reddit’s homepage visibility after Epstein’s death functions as a compositional filter that reshapes r/conspiracy’s user base, participation patterns, and linguistic integration. It focuses on whether users who arrive during this visibility event behave and write differently from those who discover the community through more organic pathways, and whether these differences persist over time. By combining behavioral, lexical, and semantic measures, the study links discovery context to toxicity, retention, and linguistic alignment.
The analysis addresses three interconnected research questions:
@@
For RQ2, we construct four mutually exclusive user cohorts based on temporal boundaries and engagement focus, and we use Kaplan-Meier survival analysis to model retention trajectories and compare users who discover the community through homepage exposure versus organic pathways. Users whose first post is about Epstein show higher retention compared to users whose first engagement is not related to Epstein, thus indicating that the Epstein suicide topic possibly attracted individuals with a stronger interest in conspiracy theories.
Finally, for RQ3, we employ Sentence-BERT embeddings to compute semantic similarity between newcomer posts and established community discourse, and we track whether users who arrive via homepage visibility exhibit persistent linguistic divergence from community norms. Users who join during the arrest have significantly higher linguistic similarity to the core users of the community. This signal remains strong across all the observation windows, while the similarity of users joining after Epstein’s death is lower across the study period. These findings suggest that users who organically discover the fringe community have a stronger language adaptation to long-term conspiracy theorists, potentially signaling a stronger intrinsic interest in the community.
@@ 2. Related work
Diffusion of conspiracy theories on social media platforms. Research shows that social media provides fertile ground for conspiratorial content to circulate at scale (Bessi et al., 2015). During moments of crisis—such as the COVID-19 pandemic—these dynamics intensify, and conspiracy narratives often spread faster than verified information (Kauk et al., 2021). Propagation has been modeled using epidemiological frameworks, revealing higher virality and persistence relative to factual news. Other work highlights cross-platform reinforcement, demonstrating that discourse emerging outside conspiracy spaces can seed and accelerate conspiracy narratives elsewhere (Paudel et al., 2021). Together, these studies emphasize the need to understand how platform mechanisms facilitate the growth and reach of conspiracy content.
@@
Detection of conspiracy theories online. Several studies have focused on detecting online conspiracy theories. combining natural-language processing, network analysis, and graph-based models. Recent studies showed that both supervised classifiers and large language models can detect conspiratorial messages with high accuracy, reaching high classification scores even in real settings (Pustet et al., 2024; Corso et al., 2025a). Another approach leverages “event-relation graphs,” mapping how events in a text are causally or temporally related; this helps flag articles whose narrative structure reveals irrational or conspiratorial linkages (Lei and Huang, 2023). Some frameworks further enrich detection by combining sentiment and topic modeling with qualitative coding, useful for uncovering subtle, implicit conspiratorial themes that keyword- or formula-based systems often miss (Langguth et al., 2023).
@@
Conspiracy engagement and user pathways. Work on conspiracy engagement shows that users rarely arrive in conspiracy communities by accident. Klein et al. (2019) trace Reddit users before their first post in r/conspiracy and show that future participants already display distinctive linguistic patterns and spend time in related subreddits. Phadke et al. (2022) follow users after they join r/conspiracy and identify multiple engagement trajectories, from stable high participation to rapid decline, linking persistent involvement to insider language and monological worldviews. Similarly, Corso et al. (2025b) analyze the linguistic traits of users active in r/conspiracy in mainstream discourse, finding unique community-specific linguistic fingerprints that identify conspiracy users in the non-fringe communities, which hold strong predictive power even years before the user’s first activity on r/conspiracy. Together, these studies treat conspiracy engagement as a selective process shaped by prior dispositions and internal adaptation, but they focus on trajectories within the community rather than on how algorithmic visibility changes who arrives in the first place.
@@
Linguistic adaptation and integration. Another line of research studies how users adapt their language to community norms. Danescu-Niculescu-Mizil et al. (2013) propose a lifecycle model in which users first align strongly with local norms and then enter a conservative phase where their style stabilizes and may drift from an evolving baseline. Studies of conspiracy discourse apply similar ideas to specific events and platforms. Schatto-Eckrodt et al. (2024) show how conspiracy narratives about Epstein’s death build on older myths and political tensions, while work on r/conspiracy documents that dramatic events bring sudden influxes of new users and shift topics and tone (Samory and Mitra, 2018; de Wildt and Aupers, 2024). These studies demonstrate that external shocks matter for discourse, but they do not compare linguistic integration across cohorts defined by how users discover the community.
@@
Platform interventions, visibility, and conspiracy communities. Several work examines how platforms intervene in problematic communities. Research on bans and deplatforming shows that removing communities does not always reduce their activity. Monti et al. (2023) find that QAnon users banned from Reddit migrate to alternative platforms and rebuild their networks, while Trujillo and Cresci (2022) show that quarantining r/The_Donald briefly reduces activity and toxicity before both rebound. Other work documents spillover effects when users carry norms and content to fringe platforms after restrictions (Horta Ribeiro et al., 2023; Russo et al., 2023). This literature treats visibility mostly as something platforms reduce through sanctions and moderation tools.
Overall, these different works show that conspiracy engagement is selective, that users adapt linguistically to community norms, and that platform decisions can reshape participation. However, they leave open how promotional algorithmic visibility, such as homepage featuring, acts as a compositional filter that changes who appears in a community and how these newcomers behave and write once they arrive. This paper addresses that gap by comparing cohorts of r/conspiracy users defined by their discovery context (organic arrest-period arrivals vs. algorithmic-visibility death-period arrivals) and by linking this compositional difference to three concrete outcomes: toxicity trajectories among core members, retention patterns for newcomers, and sentence-level semantic distance between newcomers and the established community.
3. Methods
@@
We analyze all submissions and comments posted on r/conspiracy between January 1 and December 31 2019, collected from monthly Pushshift archives of Reddit (Baumgartner et al., 2020). After filtering out deleted authors and empty or deleted texts, the dataset contains 73 32073\,320 submissions and 2 207 9052\,207\,905 comments, for a total of 2 281 2252\,281\,225 posts (Table˜1).
Table 1. Number of submissions, comments, and total posts after data cleaning.
@@
We identify Epstein-related threads with a case-insensitive regular expression over submission titles and texts that matches the complete words epstein, jeffrey epstein, ghislaine, or maxwell. We then include all comments to these submissions, which gives a focused set of posts that discuss Epstein either directly or inside those threads.
We define four mutually exclusive newcomer cohorts based on when users first post in r/conspiracy and whether that first post concerns Epstein. The cohorts are:
@@
Additionally, as shown in Figure˜2, there are two distinct spikes in public interest (proxy via Google Search trends), and we use these events as anchors to examine corresponding activity on Reddit. Overall engagement on Reddit appears largely stable; therefore, the patterns we observe are likely driven by these attention spikes, which trigger shifts within the platform. This interpretation is further supported by the fact that r/conspiracy appeared on the Reddit homepage during this period (Ellis, [n. d.]).
Figure 1. Longitudinal cohort analysis showing weekly posting activity for each of the four user cohorts.
Figure 2. Longitudinal Google Search prevalence of keywords: ‘Jeffrey Epstein’ and ‘Reddit’ during the arrest and death periods.
3.1. Toxicity time series
We study how the toxicity of core users changes around Epstein’s death by constructing a daily toxicity time series and subsequently fitting an Interrupted Time Series (ITS) model. To quantify toxicity, we compute a score between 0 and 1 for each post using Detoxify  (Hanu and Unitary, 2020), where higher values signify more offensive or hostile language. Because daily post volume varies widely, we employ a robust bootstrap procedure to obtain stable daily estimates.
@@ 3.1. Toxicity time series
We fit a segmented linear regression model with ordinary least squares (OLS), treating August 10, 2019 (the day of Epstein’s death) as the intervention point $t_{0}$ (Trujillo and Cresci, 2022). The model is:
(1) $y_{t}=\beta_{0}+\beta_{1}t+\beta_{2}D_{t}+\beta_{3}P_{t},$
@@
3.2. Lexical categories
We use Empath (Fast, 2025) to measure thematic content in posts. We focus on 17 categories that are relevant to conspiracy discourse and community dynamics, including government, politics, crime, violence, trust, conflict, secrecy, and dispute, as well as a custom paranoia category that we build from seed words such as suspicion, paranoid, watching, surveillance, threat, and conspiracy. Empath assigns each post a length-normalized score per category based on the number of category words it contains.
@@ 3.2. Lexical categories
We examine lexical shifts in two ways. First, we quantify an average within-user behavioral shift for core members (Balsamo et al., 2023). For each event (Epstein’s arrest and death), we define a one-month window before and a one-month window after the event. For every core user who posts in both windows, we compute Empath scores in the pre- and post-event months and take the difference. We then average these user-level differences by category to obtain the mean within-user shift around each event.
@@
3.3. User retention and survival analysis
@@
Using these durations and event indicators, we estimate retention with Kaplan-Meier survival curves, one per cohort. The estimator computes, at each distinct departure time, the conditional probability that users remain active given that they have not yet left, and multiplies these probabilities over time. The resulting curve shows, for each cohort, the probability that a user is still active in r/conspiracy after a given number of days since their first post. We also compute the Integrated Brier Score for each cohort using scikit-survival to summarize how well the survival model predicts observed departures over time.
3.4. Linguistic integration
To analyze linguistic integration, we measure how closely newcomers’ posts align semantically with the language of established core users over time (Danescu-Niculescu-Mizil et al., 2013). This analysis uses sentence embeddings from the all-MiniLM-L6-v2 Sentence-BERT model in the SentenceTransformers library. [^1] https://sbert.net/ The model maps each post to a 384-dimensional vector by applying the default tokenizer, truncating sequences longer than 512 tokens, and using mean pooling over token representations. We use the pre-trained model without fine-tuning, which keeps the analysis reproducible and avoids overfitting to this specific community.
@@ 3.4. Linguistic integration
(3) $\mathop{cos\_dist}(u,v)=1-\frac{u\cdot v}{\|u\|\,\|v\|}.$
@@
4. Results
4.1. Behavioral and thematic responses (RQ1)
Figure 3. Interrupted time series analysis of core users’ toxicity. The blue line shows the observed daily toxicity estimate. The red segments show the fitted pre-event and post-event trends from the OLS regression. The vertical line at $t_{0}$ marks Epstein’s death. One extreme spike above 7 standard deviations in June comes from a day with very few posts and one highly toxic message; this outlier does not drive the fitted trend because the bootstrap uses medians.
Our first research question asks how the visibility event alters the behavior of established community members. The ITS analysis in Figure˜3 shows a clear and immediate shift in how core users write. Toxicity scores drop by 1.14 standard deviations ($p<0.001$) on the day of Epstein’s death. This level change is large and negative, while the post-event slope coefficient is small and not statistically different from zero (Table˜2). In the months after the event, daily toxicity values fluctuate around a lower mean than in the pre-event period, without a strong upward rebound in the fitted trend.
Table 2. Interrupted time series regression estimates for core users’ toxicity.

### v2

# Effects of Mainstream Visibility on Conspiracy Communities: Reddit after Epstein’s ‘Suicide’

@@ Abstract
Following the death of Jeffrey Epstein, the subreddit r/conspiracy experienced a significant visibility shock that brought mainstream users into direct contact with established conspiracy narratives. In this work, we explore how large-scale surges in public attention reshape participation and discourse within online conspiracy communities. We ask whether a sudden increase in exposure changes who join r/conspiracy, how long they stay, and how they adapt linguistically, compared with users who arrive through organic discovery. Using a computational framework that combines toxicity scores, survival analysis, and lexical and semantic measures over a period of 12 months, we observe that mainstream visibility is is associated with patterns consistent with a selection mechanism rather than a simple amplifier. Users who join the conspiracy community during the arrest-period tend to show higher linguistic similarity to core users, especially regarding linguistic and thematic norms and showing more stable engagement over time. By contrast, users who arrive during the height of public visibility remain semantically distant from core discourse and participate more briefly. Overall, we find that mainstream visibility is connected with changes in audience size, community composition, and linguistic cohesion. However, incidental exposure during attention shocks does not typically produce durable, integrated community members. These results provide a more nuanced understanding of how external events and platform visibility influence the growth and evolution of conspiracy spaces, offering insights for the design of responsible and transparent recommendation systems.
Introduction
@@
Conspiracy theories are false or unverified narratives that allege secret, often sinister plots or events involving groups, organizations, or governments (Douglas and Sutton 2023). They pose significant risks since they can change individuals’ perception of real events (Uscinski et al. 2017). Social media platforms provide especially fertile ground for their circulation, in part due to the emergence of fringe online communities (Rollo et al. 2022; Spohr 2017). Such theories range from relatively benign conjectures to more harmful beliefs that amplify misinformation, heighten polarization, and potentially lead to serious real-world consequences.
@@
Online conspiracy communities are critical environments for understanding how digital platforms shape ideological discourse. Yet, it’s not clear how mainstream visibility of these fringe communities plays a role in determining who joins these communities, how long they stay, and whether they culturally integrate (Klein et al. 2019; Phadke et al. 2022; Venturini 2022; Zeng et al. 2022; Corso et al. 2025c). Existing work on conspiracy engagement shows that users follow distinct pathways into and through conspiracy spaces and that linguistic adaptation and community ties play a central role in deeper involvement (Klein et al. 2019; Phadke et al. 2022). At the same time, research on platform interventions and deplatforming shows that conspiracy communities can be unusually resilient and mobile, rebuilding networks elsewhere even after bans (Monti et al. 2023). What remains unclear is how platform-level visibility itself, before any formal intervention, shapes who enters these communities and what kinds of membership emerge.
@@
Reddit offers a useful setting for such a research question (Proferes et al. 2021). It organizes content into topical subreddits and uses ranking and recommendation systems to decide which posts appear on the site’s homepage. When a subreddit reaches the homepage, it exposes its community to millions of users who are not subscribed to it and might not otherwise encounter its content. The subreddit r/conspiracy is Reddit’s largest conspiracy community, with more than two million members and a long record of discussions that range from historical events to contemporary politics. This combination of scale, public visibility, and stable infrastructure makes r/conspiracy an ideal case for studying how visibility events affect community composition, participation, and language (de Wildt and Aupers 2024).
@@
The death of financier Jeffrey Epstein on August 10, 2019, created a sharp visibility shock (Schatto-Eckrodt et al. 2024). The peculiar circumstances of his death aligned fringe conspiracy discourse with mainstream skepticism and brought r/conspiracy and its 2 million members to the fore, reaching the platform’s homepage, as reported by users from the same community. [^1] https://reddit.com/r/conspiracy/comments/cojrq9222https://reddit.com/r/conspiracy/comments/covvj0 This visibility event changed the community’s reach without changing its formal rules or moderation structures, and it did so over a short, well-defined time window as visible in Figure˜8
@@
In this work, we examine how Reddit’s homepage visibility after Epstein’s death is associated with changes in community composition consistent with a compositional filter that reshapes r/conspiracy’s user base, participation patterns, and linguistic integration. We focus on whether users who arrive during this visibility event behave and write differently from those who discover the community through more organic pathways, and whether these differences persist over time. By combining behavioral, lexical, and semantic measures, this study links discovery context to toxicity, retention, and linguistic alignment. Specifically, we aim to answer the following research questions:
@@
For RQ2, we construct four mutually exclusive user cohorts based on temporal boundaries and engagement focus, and we use Kaplan-Meier survival analysis to model retention trajectories and compare users who enter during the heightened visibility of the death-period versus those who arrived during the earlier arrest-period. Users whose first post is about Epstein show higher retention compared to users whose first engagement is not related to Epstein, thus indicating that the Epstein suicide topic possibly attracted individuals with a stronger interest in conspiracy theories.
Finally, for RQ3, we employ Sentence-BERT embeddings to compute semantic similarity between newcomer posts and established community discourse, and we track whether users who arrive via homepage visibility exhibit persistent linguistic divergence from community norms. Users who join during the arrest have significantly higher linguistic similarity to the core users of the community. This signal remains strong across all the observation windows, while the similarity of users joining after Epstein’s death is lower across the study period. These findings suggest that users who discover the conspiracy community during the arrest period show a stronger language adaptation to long-term conspiracy theorists, potentially signaling a stronger intrinsic interest in the community.
@@
Diffusion of conspiracy theories on social media platforms. Research shows that social media provides fertile ground for conspiratorial content to circulate at scale (Bessi et al. 2015). During moments of crisis—such as the COVID-19 pandemic—these dynamics intensify, and conspiracy narratives often spread faster than verified information (Kauk et al. 2021). Propagation has been modeled using epidemiological frameworks, revealing higher virality and persistence relative to factual news. Other work highlights cross-platform reinforcement, demonstrating that discourse emerging outside conspiracy spaces can seed and accelerate conspiracy narratives elsewhere (Paudel et al. 2021). Together, these studies emphasize the need to understand how platform mechanisms facilitate the growth and reach of conspiracy content.
@@
Detection of conspiracy theories online. Several studies have focused on detecting online conspiracy theories. combining natural-language processing, network analysis, and graph-based models. Recent studies showed that both supervised classifiers and large language models can detect conspiratorial messages with high accuracy, reaching high classification scores even in real settings (Pustet et al. 2024; Corso et al. 2025a). Recent work also showed how LLMs also can be used as proxy for simulating conspiracy-inclined individuals (Corso et al. 2025b). Another approach leverages “event-relation graphs,” mapping how events in a text are causally or temporally related; this helps flag articles whose narrative structure reveals irrational or conspiratorial linkages (Lei and Huang 2023). Some frameworks further enrich detection by combining sentiment and topic modeling with qualitative coding, useful for uncovering subtle, implicit conspiratorial themes that keyword- or formula-based systems often miss (Langguth et al. 2023).
@@
Conspiracy engagement and user pathways. Work on conspiracy engagement shows that users rarely arrive in conspiracy communities by accident. Klein et al. (2019) trace Reddit users before their first post in r/conspiracy and show that future participants already display distinctive linguistic patterns and spend time in related subreddits. Phadke et al. (2022) follow users after they join r/conspiracy and identify multiple engagement trajectories, from stable high participation to rapid decline, linking persistent involvement to insider language and monological worldviews. Similarly, Corso et al. (2025c) analyze the linguistic traits of users active in r/conspiracy in mainstream discourse, finding unique community-specific linguistic fingerprints that identify conspiracy users in the non-fringe communities, which hold strong predictive power even years before the user’s first activity on r/conspiracy. Together, these studies treat conspiracy engagement as a selective process shaped by prior dispositions and internal adaptation, but they focus on trajectories within the community rather than on how mainstream visibility changes who arrives in the first place.
@@
Linguistic adaptation and integration. Another line of research studies how users adapt their language to community norms. Danescu-Niculescu-Mizil et al. (2013) propose a lifecycle model in which users first align strongly with local norms and then enter a conservative phase where their style stabilizes and may drift from an evolving baseline. Studies of conspiracy discourse apply similar ideas to specific events and platforms. Schatto-Eckrodt et al. (2024) show how conspiracy narratives about Epstein’s death build on older myths and political tensions, while work on r/conspiracy documents that dramatic events bring sudden influxes of new users and shift topics and tone (Samory and Mitra 2018; de Wildt and Aupers 2024). These studies demonstrate that external shocks matter for discourse, but they do not compare linguistic integration across cohorts defined by how users discover the community.
@@
Platform interventions, visibility, and conspiracy communities. Several work examines how platforms intervene in problematic communities. Research on bans and deplatforming shows that removing communities does not always reduce their activity. Monti et al. (2023) find that QAnon users banned from Reddit migrate to alternative platforms and rebuild their networks, while Trujillo and Cresci (2022) show that quarantining r/The_Donald briefly reduces activity and toxicity before both rebound. Other work documents spillover effects when users carry norms and content to fringe platforms after restrictions (Horta Ribeiro et al. 2023; Russo et al. 2023). This literature treats visibility principally as something platforms reduce through sanctions and moderation tools.
Overall, these different works show that conspiracy engagement is selective, that users adapt linguistically to community norms, and that platform decisions can reshape participation. However, the role of sudden visibility surges in filtering community membership and reshaping discourse remains under-explored. We address this by comparing newcomer cohorts arriving during two distinct attention spikes—the arrest-period and the high-visibility death-period—linking these arrival windows to core-user toxicity, newcomer retention, and linguistic integration to evaluate how large-scale interest shocks influence community dynamics.
Methods
@@
We analyze all submissions and comments posted on r/conspiracy between January 1 and December 31 2019, collected from monthly Pushshift archives of Reddit (Baumgartner et al. 2020). After filtering out deleted authors and empty or deleted texts, the dataset contains 73 32073\,320 submissions and 2 207 9052\,207\,905 comments, for a total of 2 281 2252\,281\,225 posts (Table˜1).
@@
Table 1: Number of submissions, comments, and total posts after data cleaning.
We identify Epstein-related threads with a keyword-based matching approach over submission titles and texts on the following set of keywords: epstein, jeffrey epstein, ghislaine, or maxwell. We then include all comments to these submissions, which gives a focused set of posts that discuss Epstein either directly or inside those threads. We chose the previously mentioned keywords due their strict connection with our case study, which helps us in maximizing the precision of the matching, even at the cost of a lower recall, to ensure the "engaged" cohorts are truly topical. We define four mutually exclusive newcomer cohorts based on when users first post in r/conspiracy and whether that first post concerns Epstein. The cohorts are:
@@
Additionally, as shown in Figure˜7 in the Appendix, there are two distinct spikes in public interest (proxy via Google Search trends), and we use these events as anchors to examine corresponding activity on Reddit. Overall engagement on Reddit appears largely stable; therefore, the patterns we observe are likely driven by these attention spikes, which trigger shifts within the platform.
Figure 1: Longitudinal cohort analysis showing weekly posting activity for each of the four user cohorts.
Toxicity time series
We study how the toxicity of core users changes around Epstein’s death by constructing a daily toxicity time series and subsequently fitting an Interrupted Time Series (ITS) model. To quantify toxicity, we compute a score between 0 and 1 for each post using Detoxify  (Hanu and Unitary 2020), where higher values signify more offensive or hostile language. Because daily post volume varies widely, we employ a robust bootstrap procedure to obtain stable daily estimates.
@@
We fit a segmented linear regression model with ordinary least squares (OLS), treating August 10, 2019 (the day of Epstein’s death) as the intervention point $t_{0}$ (Trujillo and Cresci 2022). The model is:
$y_{t}=\beta_{0}+\beta_{1}t+\beta_{2}D_{t}+\beta_{3}P_{t},$ (1)
@@
Lexical categories
We use Empath (Fast 2025) to measure thematic content in posts. We focus on 17 categories that are relevant to conspiracy discourse and community dynamics, including government, politics, crime, violence, trust, conflict, secrecy, and dispute, as well as a custom paranoia category that we build from seed words such as suspicion, paranoid, watching, surveillance, threat, and conspiracy. Empath assigns each post a length-normalized score per category based on the number of category words it contains.
@@
We examine lexical shifts in two ways. First, we quantify an average within-user behavioral shift for core members (Balsamo et al. 2023). For each event (Epstein’s arrest and death), we define a one-month window before and a one-month window after the event. For every core user who posts in both windows, we compute Empath scores in the pre- and post-event months and take the difference. We then average these user-level differences by category to obtain the mean within-user shift around each event.
@@
User retention and survival analysis
@@
Using these durations and event indicators, we estimate retention with Kaplan-Meier survival curves, one per cohort. The estimator computes, at each distinct departure time, the conditional probability that users remain active given that they have not yet left, and multiplies these probabilities over time. The resulting curve shows, for each cohort, the probability that a user is still active in r/conspiracy after a given number of days since their first post. We also compute the Integrated Brier Score for each cohort to summarize how well the survival model predicts observed departures over time. Lower values of IBS indicate that the timing of user churn within a cohort is more regular and closely aligns with the model’s survival estimates, whereas higher scores reflect more diverse participation patterns.
Linguistic integration
To analyze linguistic integration, we measure how closely newcomers’ posts align semantically with the language of established core users over time (Danescu-Niculescu-Mizil et al. 2013). This analysis uses sentence embeddings from the all-MiniLM-L6-v2 Sentence-BERT model in the SentenceTransformers library. [^3] https://sbert.net The model maps each post to a 384-dimensional vector by applying the default tokenizer, truncating sequences longer than 512 tokens, and using mean pooling over token representations. We use the pre-trained model without fine-tuning, which keeps the analysis reproducible and avoids overfitting to this specific community.
@@
$\mathop{cos\_dist}(u,v)=1-\frac{u\cdot v}{\|u\|\,\|v\|}.$ (3)
@@
Results
Behavioral and thematic responses (RQ1)
Figure 2: Interrupted time series analysis of core users’ toxicity. The blue line shows the observed daily toxicity estimate. The red segments show the fitted pre-event and post-event trends from the OLS regression. The vertical line at $t_{0}$ marks Epstein’s death. One extreme spike above 7 standard deviations in June comes from a day with very few posts and one highly toxic message; this outlier does not drive the fitted trend because the bootstrap uses medians.
Our first research question asks how the visibility event alters the behavior of established community members. The ITS analysis in Figure˜2 shows a clear and immediate shift in how core users write. Toxicity scores drop by 1.14 standard deviations ($p<0.001$) on the day of Epstein’s death. This level change is large and negative, while the post-event slope coefficient is small and not statistically different from zero (Table˜2). In the months after the event, daily toxicity values fluctuate around a lower mean than in the pre-event period, without a strong upward rebound in the fitted trend.